OpenStudy (anonymous):

How tall is a cell phone tower that casts a 120-foot shadow when the sun is at a 83 degree of elevation? (round to nearest foot)

OpenStudy (amistre64):

tan(83) = h/d d tan(83) = h 120 tan(83) = h

OpenStudy (anonymous):

\[tan(83)=\frac{h}{120}\] \[h=120tan(83)\]

OpenStudy (anonymous):

i am too slow

OpenStudy (amistre64):

lol...its that equation editor; eats alot of time :)

OpenStudy (anonymous):

using latex. don't like editor plus i get latex practice.

OpenStudy (amistre64):

i get to many typos to do latex :)

OpenStudy (anonymous):

that is why i am doing it. increase my speed plus memorize commands.